### v1

# RAG-Driver: Generalisable Driving Explanations with Retrieval-Augmented In-Context Learning in Multi-Modal Large Language Model

Abstract
Robots powered by ‘blackbox’ models need to provide human-understandable explanations which we can trust. Hence, explainability plays a critical role in trustworthy autonomous decision-making to foster transparency and acceptance among end users, especially in complex autonomous driving. Recent advancements in Multi-Modal Large Language models (MLLMs) have shown promising potential in enhancing the explainability as a driving agent by producing control predictions along with natural language explanations. However, severe data scarcity due to expensive annotation costs and significant domain gaps between different datasets makes the development of a robust and generalisable system an extremely challenging task. Moreover, the prohibitively expensive training requirements of MLLM and the unsolved problem of catastrophic forgetting further limit their generalisability post-deployment. To address these challenges, we present RAG-Driver, a novel retrieval-augmented multi-modal large language model that leverages in-context learning for high-performance, explainable, and generalisable autonomous driving. By grounding in retrieved expert demonstration, we empirically validate that RAG-Driver achieves state-of-the-art performance in producing driving action explanations, justifications, and control signal prediction. More importantly, it exhibits exceptional zero-shot generalisation capabilities to unseen environments without further training endeavours [^1] Project Page: https://yuanjianhao508.github.io/RAG-Driver.
Index Terms:
Autonomous driving, multi-modal language model, end-to-end driving, domain generalisation
I Introduction
Driven by the emerging development of deep learning, autonomous driving has observed a paradigm shift from rules-based decision systems [66, 21] to data-driven learning-based approaches [28, 6, 36]. However, this comes at the cost of transparency in decision-making, especially for end-to-end autonomous driving systems which are considered black-box in nature [13]. Thus, in addition to precision in action control, explanation provision is key in ensuring trustworthy decision-making to reconcile the system’s decisions with end-user expectations to foster confidence and acceptance [79, 8, 57] in dynamic driving environments.
Traditional approaches have mainly relied on attention visualisation [5, 7, 55] as a proxy to rationalise the decisions of the black-box systems or auxiliary intermediate tasks such as semantic segmentation [25, 32], object detection [16, 31], and affordance prediction [68, 45] provide meaningful intermediate representation for decision-making. However, these methods do not engage end-users in the dialogue as they are one-directional and not readily comprehensible by the general users for the purpose of fostering trust and confidence. An alternative promising approach is the integration of natural language explanations [38, 33, 54], in particular through Multi-Modal Large Language Models (MLLMs) [1, 70]. These models, pre-trained on extensive web-scale datasets, demonstrate remarkable reasoning capacity, enabling the transformation of complex vehicular decision-making processes into more understandable narrative formats, thereby offering a new layer of explainability to conventional systems.
While several early attempts have demonstrated the potential of MLLMs as general explainable driving agents [78, 76, 51], these methods fall short of human-level understanding. One of the limitations is their failure to generalise to unseen environments. A primary obstacle is the lack of high-quality annotated data [56], coupled with the significant domain shift across various datasets [23], which hinders the models’ generalisation capacity to novel environments outside of the training data distribution. Another critical challenge is the prohibitively expensive training requirement and the unsolved problem of catastrophic forgetting [39], which make re-training or fine-tuning impractical solutions due to the immense computational demands and severe performance degradation. Consequently, this further limits the models’ generalisability after deployment, as they struggle to effectively utilise new data in constantly evolving environments and driving scenarios.
To address these challenges, we introduce RAG-Driver, a novel retrieval-augmented multi-modal large language model tailored for generalisable and explainable end-to-end driving. As illustrated in Fig.1, it outputs natural language texts corresponding to (1) the driving action and (2) justification of that driving action along with (3) numerical control signals based on driving videos. The natural language texts are aligned with control signals during in-context learning [10] to enable faithful introspective explanation provision. The novelty of RAG-Driver is the integration of retrieval-augmented in-context learning (RA-ICL) mechanisms that significantly improve generalisation performance in unseen driving environments. It allows efficient recall of similar driving scenarios as contextual information augmenting MLLM prediction through implicit meta-optimisation (Sec. III-C). Through extensive experiments, we show that RAG-Driver outperforms existing methods on both in-domain deployments as well as deployment in unseen environments (without any fine-tuning). By being grounded in analogical demonstrations, our framework significantly reduces the need for continuous retraining while enhancing the generalisability and quality of generated explanatory texts. Our primary contributions are as follows:
  1. 1.
    Proposing a novel retrieval-augmented in-context learning method for Multi-Modal Large Language Model (MLLM) based generalisable and explainable driving.
  2. 2.
    Achieving state-of-the-art introspective driving explanation performance on the standard benchmark BDD-X [38].
  3. 3.
    Demonstrating exceptional zero-shot generalisation capacity to unseen scenarios without training effort through a customised dataset Spoken-SAX featuring video sequences narrated by a professional driving instructor.
II Related Work
II-A Explainable End-to-End Autonomous Driving
End-to-end learned driving [13] maps directly from raw sensor input to vehicle control signals. This data-driven, joint optimisation of perception, prediction, and planning can be simple and efficient [28]. In this area, various learning-based approaches, including behavior cloning [6, 12, 61, 83, 60], inverse optimal control [81, 65, 75], and reinforcement learning [36, 71, 11, 44] are promising. A key area of focus in this area is explainability [80], which is crucial for improving transparency and building trust towards wider public acceptance of autonomous systems [54, 22, 57]. One line of works leverages attention visualisation – either to directly identify salient regions of input images that are important for driving decision-making [37, 5, 7] or to assist feature aggregation for downstream motion planning tasks [55, 15, 77, 63]. Another line of work uses intermediate auxiliary tasks such as semantic segmentation [25, 32], object detection [16, 31], and affordance prediction [68, 45] which help to decoder latent representations to human-understandable representation. While these methods provide explainable mechanisms by associating decision-making processes with semantic or visual representations, they are not readily comprehensible by the general users for the purpose of fostering trust and confidence.
Alternatively, recent research shows promise in utilising natural language explanation. Several works develop specialist explainers [38, 41] using attention alignment in visual input and textual generation for grounded driving action explanation. ADAPT [33] uses a vision-language transformer with separate decoder for caption generation alongside control signal prediction. More recently, several works explore the potential of Multi-modal Large Language Models (MLLMs). Works such as, DriveGPT4 [78], Lingo [54], and DrivingMLM [76] have shown promising potential in general question-answering for driving and action planning. However, a common obstacle in both specialist and MLLM-based generalist models is the scarcity of data due to expensive annotation costs and significant domain gaps between different datasets, making the development of a robust and generalisable model an extremely challenging task. In our work, we aim to overcome these obstacles by employing a more robust inference paradigm of retrieval-augmented in-context learning to bridge domain gaps and circumvent the need for annotations in new domains.
Figure 1: RAG-Driver Overview: Given a query comprising a video of the current driving scenario and its corresponding control signal, the process starts with the input being fed into a Retrieval Engine. It searches within a memory database for driving experiences that are similar to the current scenario, thereby providing relevant in-context learning samples. Subsequently, a Multi-Modal Large Language Model (MLLM) processes both the current query and the retrieved in-context learning samples. Based on high-level task instructions, the model engages in various prediction tasks: Action Explanation, Action Justification, and Next Control Signal Prediction.
II-B Multi-Modal Large Language Model
Recent advancements in Large Language Models (LLMs) have paved the way for the emergence of Multi-modal Large Language Models (MLLMs) [1, 70]. Benefiting from scalable transformer-based architecture and web-scale training data, these models have demonstrated notable capabilities in general-purpose visual understanding tasks. One line of work focuses on modality fusion in the latent space, offering a scalable, end-to-end solution for MLLMs. For instance, Flamingo [3] and BLIP2 [43] fuse visual tokens to frozen LLM through gated attention and query transformers, respectively. LLaVA [48] and MiniGPT4 [86] use simple Multi-Layer Perceptron (MLP) with visual instruction tuning to align the pre-trained image encoder to LLM. Most relevant to us is the line of work focusing on video-language models such as Video-LLaVA [46] and Video-LLaMA [82], which integrate pre-trained video encoders into LLMs using strategies similar to those in image-based models.
With remarkable perception and reasoning capacity, MLLMs reveal promising potential in various robotics tasks such as reasoning [20, 2, 30] and planning [62, 9, 47, 35]. The most similar to us is the idea that engages a generalist foundation model for end-to-end embodied agent. PaLM-e [20] injects images, state estimates, and other sensor modalities to LLM and autoregressively produces natural language commands. RT-2 [9] and RT-X [58] fine-tune on image and low-level robot control signal pairs to perform end-to-end robot control. Specifically in driving, numerous approaches leverage language-only LLM for decision-making [79] and then augment it with external perception module feedback [53, 24, 14], designed chain-of-thought reasoning template [53, 52] or downstream planner [67, 69, 14] to form a system-level driving agent. Another more relevant line of work is end-to-end driving agents. DriveGPT4 [78] leverages a finetuned video-language model Valley [50] on driving specific visual instruction tuning based on BDD-X [38]. Dolphins [51] further use the designed Grounded chain-of-thought to enhance reasoning capacity. DrivingMLM [76] and Reason2Drive [56] scale up the driving visual instruction tuning dataset through simulator-based data engine and annotation of existing large-scale datasets, respectively. While these approaches have demonstrated the potentials of MLLM, the prohibitively expensive training cost and the unsolved problem of catastrophic forgetting, which makes re-training or fine-tuning post-deployment challenging, further limit their generalisation capacity to unseen driving environments. To solve this problem, we leverage a training-free retrieval-augmented in-context learning mechanism.
II-C In-Context Learning and Retrieval-Augmented Generation
While LLMs demonstrate strong generative and reasoning capacity, there are still several issues associated with their output, such as hallucination [29], and slow knowledge updates [34]. In-context Learning (ICL) [10, 18] has emerged as a promising approach in LLM inference, potentially addressing several of these issues. This paradigm involves providing a test query alongside a few demonstration examples as contextual information. The LLM then generates an output for the test instance based on analogies drawn from context, without any updates to its parameters [64]. While ICL has been observed to enhance generalisability in various Natural Language Processing (NLP) tasks, its application in multi-modal contexts remains less explored, potentially due to the challenges associated with curating structured high-quality multi-modal ICL datasets. Retrieval-Augmented Generation (RAG) [42] is another important inference paradigm for LLM. It provides an external knowledge database to augment the compressed model knowledge within LLM in inference by dynamically retrieving relevant information pieces as contextual information. One of its promising applications is leveraging a systematic approach to curate In-Context Learning (ICL) examples. In this work, we build upon these inference paradigms and extend their application to Multimodal Large Language Models (MLLMs). We introduce a retrieval-augmented in-context learning mechanism through a curated multi-modal driving in-context instruction tuning dataset and a vector similarity-based retrieval engine specifically tailored for driving applications.
III Method
RAG-Driver is a retrieval-augmented, multi-modal large language model (MLLM) for generalisable explainable end-to-end driving. Its multi-tasking capabilities encompass three key areas: (1) Action Explanation, providing a human-understandable driving action description; (2) Action Justification, elucidating the rationale behind specific driving actions; and (3) Next Control Signal Prediction, forecasting upcoming control signals in response to the driving conditions. As shown in Fig. 1, it is composed of two primary components: a unified perception and planning unit built upon an MLLM backbone and a memory unit built upon a hybrid vector and textual database. These components interact through a retrieval engine, enabling robust multi-modal in-context learning (ICL) during decision-making.
III-A Multi-modal Large Language Model Architecture
Following the successful MLLM paradigm of Video-LLaVA [46], we align visual and language embedding through visual instruction tuning. We leverage a pre-trained video encoder and LLM and then inject the video embedding into LLM through a MLP projector to build a fully differentiable MLLM.
Video Encoder We adopt the pre-trained LanguageBind video encoder [85] as our frozen visual backbone $f_{v}$, which is based on a ViT-B/32 vision transformer [19]. As shown in Fig. 2, given an input video frame sequence $V_{i}=\left\{v_{i}^{1},v_{i}^{2},\ldots,v_{i}^{k}\right\}\in\mathbb{R}^{3\times k\times 224\times 224}$, we first split the video into multiple temporal sequences, each consisting of patches that share the same spatial location across different frames. These patches are then transformed through a linear projection for a vision transformer to output video embedding $\bm{z}_{vo}\in\mathbb{R}^{2048\times 1024}$. The video encoder is pre-trained with video-language contrastive learning (i.e. CLIP4clip [49]) without further fine-tuning.
Figure 2: Video Encoder architecture. Video is first split into $k\times 32\times 32$ patches concatenated in time, where these patches are linear projected to video embedding. Then, the model is trained with video-language contrastive learning (CLIP4clip) [49] to obtain language-align video representation.
Cross-Modality Projector We then leverage a two-layer MLP to project and align the encoded video embedding $\bm{z}_{vo}$ with language token embeddings $\bm{z}_{v}\in\mathbb{R}^{2048\times 4096}$.
$f_{p}(\bm{z}_{vo})=\text{GELU}\left(W_{2}\cdot\text{GELU}\left(W_{1}\cdot\bm{z}_{vo}\right)\right)$ (1)
In particular, the projector $f_{p}$ takes form in Eq. 1, where we use GELU [26] as an activation function. We train the projector with a two-stage training strategy as detailed in Sec. III-B.
Large Language Model Backbone Finally, the LLM takes in both aligned video embedding $\bm{z}_{v}$ and language embedding $\bm{z}_{t}$ of textual context information and task instruction to predict both textual action explanation and numerical control signal. We adopt Vicuna 1.5 7B [84], which is instruction-tuned based on LLaMA2 [72] as our LLM backbone. For decoder-only LLM conditioned on the multi-modal contextual prefix $\bm{z}_{1:n}=[\bm{z}_{v},\bm{z}_{t}]$ of length $N$, the joint probability of output $\bm{x}_{n+1:L}$ as in Eq. 2, where $P_{\theta}$ is transformer-based LLM backbone parameterized by $\theta$.
$P(\bm{x}_{n+1:L}\mid\bm{z}_{1:n})=\prod_{l=n+1}^{L}P_{\theta}(\bm{x}_{l}\mid\bm{x}_{1:l-1},\bm{z}_{1:n})$ (2)
Each output token $\bm{x}_{l}$ is then sampled auto-regressively based on previous output and context and finally decode to language space through a text de-tokenizer.
III-B Training Strategy
Following the visual instruction tuning paradigm [48, 46], we employ a two-stage training strategy to progressively enable cross-modality alignment and multi-task driving capacity. In both stages, we leverage the same next-token prediction cross-entropy loss as in Eq. 3 aiming to maximize the conditional answer likelihood in Eq. 2, where $\bm{y}$ is the ground truth token.
$\mathcal{L}_{CE}=-\sum_{i=n+1}^{L}\bm{y}_{l}\log P(\bm{x}_{l}\mid\bm{z}_{1:n})$ (3)
Pre-training In the first pre-training stage, we only train the cross-modality projector while freezing the visual encoder and LLM. We use a subset VIDAL-10M [85] which contains 3 million video-caption pairs. This achieves an alignment between visual and language features by projecting the pre-trained video embeddings into language tokens understandable by LLM.
Supervised In-context Instruction Tuning While state-of-the-art LLM exhibits zero-shot ICL capability, several works [18] and our ablation study Sec. IV-D have shown further improvement when specifically trained on curated ICL demonstrations. In particular, we construct a multi-modal instruction tuning dataset with structured ICL examples based on the BDD-X dataset [38] resulting in 16K video question and answer pairs. As shown in Fig. 3, for an 8-frame driving video sequence with associated control signals—speed, course, acceleration, and curvature as the current query, we use the retrieval mechanism as in Sec. III-C to retrieve 2 relevant driving experiences, which are then prefixed to the current query as ICL examples. The dataset is tailored to support three distinct tasks, each represented through question-answer pairs in natural language. Note that, with (1) Action Explanation and (2) Justification naturally represented as natural language, (3) Control Signal prediction is also formed as language token prediction; this is feasible due to the distinct mapping of numerical values to specific tokens within the language model dictionary.
Figure 3: Query Sample. A single query includes three primary components: (1) Contextual System Information, a constant element across all queries that offers high-level task instruction; (2) Current Query, comprising all driving information and specific task instruction in natural language question-answer format, except driving video (i.e. $<$video$>$) as visual input. Note that the red sentence is the expected output from MLLM; (3) Multiple In-Context Learning Examples, retrieved from memory, which serve as an analogical demonstration of a complete reasoning process—from sensor input in driving scenarios to action justification and control prediction.
III-C Retrival-Augmented In-context Learning
Another critical component of the system is the memory unit, which consists of a database and a retrieval engine. The database incorporates vectorised video embedding $\bm{z}_{vo}$, which is extracted with the same video encoder as in Sec. III-A and control signals $\bm{c}\in\mathbb{R}^{28}$ directly from sensor recording. Each vector is uniquely associated with the corresponding human expert textual explanation and justification from train samples as in Sec. III-B.
Retrieval Mechanism To perform the retrieval, we first leverage a lightweight MLP projector of the same structure as in Eq. 1 to project the heterogeneous video and control signal embedding into the same hybrid embedding $\bm{s}\in\mathbb{R}^{1024}$ through metric learning [27]. In particular, we adopt a triplet loss with Euclidean distance as shown in Eq. 4:
$\mathcal{L}_{Tri}(\bm{a},\bm{p},\bm{n})=\operatorname*{max}(||\bm{a}-\bm{p}||_{2}-||\bm{a}-\bm{n}||_{2}+\operatorname{margin},0)$ (4)
where the positive $(\bm{a},\bm{p})$ and negative pairs $(\bm{n},\bm{p})$ between hybrid embedding $\bm{s}$ are selected based on text similarity (i.e., TF-IDF Score) of driving action and justification in the BDD-X training set as we aim to form the metric space such that the scenarios lead to similar driving actions close together and vice versa. This approach addresses the limitations of relying solely on visual similarity, which we have empirically found can result in sub-optimal performance (Sec. IV-D), and also solves the problem of heterogeneous sensor inputs that are hard for similarity comparison. We then perform retrieval through an efficient vector similarity search. Given a query vector $\bm{s}^{q}$ the cosine similarity between the query vector and each vector in the database is computed as follows:
$S_{c}(\bm{s}_{q},\bm{s}^{(i)})=\frac{\bm{s}_{q}\cdot\bm{s}^{(i)}}{\|\bm{s}_{q}\|\|\bm{s}^{(i)}\|}$ (5)
Subsequently, we consistently select the two most relevant driving samples based on this similarity score. These samples represent the entire reasoning process, from contextual information to question-answer pairs, as illustrated in Fig. 3.
Retrieval Augmented In-Context Learning (RA-ICL) To perform the RA-ICL, we prefix retrieved samples before the current query, facilitating an implicit gradient descent through the meta-optimiser of LLM, as proved in [17]. This approach is also applicable to MLLM with architecture specified in Sec. III-A. For a given transformer-based pre-trained MLLM modeling the prefix-conditioned output probability as in Eq. 2, consider one head of the multi-head attention in a single transformer block as follows:
$\mathcal{F}_{\mathrm{ICL}}$ $(\bm{z}_{1:n})=\operatorname{Attention}(Q,K,V)$ (6)
$=$ $W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\operatorname{Softmax}\left(\frac{\left(W_{K}\bm{z}_{1:n}\right)^{T}(W_{Q}\bm{z}_{1:n})}{\sqrt{d^{i}}}\right)$
where $\bm{z}_{1:n}=\left[\bm{z}_{icl};\bm{z}_{q}\right]$ represented the conditional prefix consist with ICL example $\bm{z}_{icl}$ and current query embeddings $\bm{z}_{q}$, respectively. $W_{Q,K,V}\in\mathbb{R}^{d^{i}\times d^{o}}$ is the linear transformation on query, key, and value in attention block, respectively. Now, in [17] in-context learning was shown to effectively achieve a meta-optimisation to implicitly update the model with an estimated meta-gradient. We provide in our work a new, alternative derivation of this, supported by more recent work in [40] on the following softmax-free linear attention expressions
$\mathcal{F}_{\mathrm{ICL}}(\bm{z}_{1:n})=W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\left(W_{K}\left[\bm{z}_{icl};\bm{z}_{q}\right]\right)^{T}W_{Q}\bm{z}_{1:n}$ (7)
This can be simplified, with a more detailed derivation in Sec. -D, as
$\mathcal{F}_{\mathrm{ICL}}(\bm{z}_{1:n})=(\Delta{}W_{ICL}+W_{ZSL})W_{Q}\bm{z}_{1:n}$ (8)
where we separate out the terms $W_{ZSL}$ independent of the ICL examples and dependent solely on the current query from those $\Delta{}W_{ICL}$ dependent on the ICL examples, given by:
$\Delta{}W_{ICL}$ $=\sum_{i}~{}W_{V}\bm{z}_{icl,i}(W_{K}\bm{z}_{icl,i})^{T}$ (9)
$W_{ZSL}$ $=W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T}$
Now, with more detail in Sec. -D, a forward-pass
$\mathcal{F}(\mathbf{x})=(W_{0}+\Delta{}W)\mathbf{x}$ (10)
through a linear layer $\mathcal{F}$ after its weights $W_{0}$ have been updated by $\Delta{}W$ has the weight updates coupled to the input in the form
$\Delta{}W\mathbf{x}=\sum_{i}\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$ (11)
where $\mathbf{x}_{i},\mathbf{y}_{i}$ are the (mini-batch) input and output to the layer that resulted in the weight update when backpropagating loss $L$ and with learning rate $\eta$. Therefore we have a weighted sum of dot products, which is akin to an attention mechanism. Indeed, by inspection of similar dot-product expressions $W_{V}\bm{z}_{icl,i}(W_{K}\bm{z}_{icl,i})^{T}\leftrightarrow\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$ we note that we match the form for the linear layer above
$(W_{ZSL}+\Delta{}W_{ICL})W_{Q}\bm{z}_{1:n}\leftrightarrow(W+\Delta{}W)\mathbf{x}$
This can therefore be interpreted to say that the output of the attention is adjusted in a meta-optimal way to conform to the samples provided as input context, much like gradient descent on the linear layer would adjust that layer to conform to the mini-batch training data, but crucially in the case of RAG-Driver, without backpropagation.
RA-ICL serves as an efficient inference method boosting the performance of MLLM in explainable driving without further training effort, where we empirically verify it is extremely effective in boosting the model prediction performance and generalisation capacity.
IV Experiments
IV-A Settings and Datasets
We empirically evaluate the proposed Retrieval-augmented In-Context Learning (RA-ICL) framework within the Multi-modal Large Language Model (MLLM), targeting explainable driving applications. We aim to validate its efficacy in general driving scenarios with a focus on two main aspects: (1) explainability in driving action explanation and justification. (2) Control Signal Prediction. We conduct experiments with the BDD-X [38] dataset, which is a widely adopted benchmark in explainable driving, comprising 77-hour videos across the US under different road and weather conditions. We customize the format as shown in Fig. 3, resulting in 16,803 and 2,123 video question-answering pairs for training and testing, respectively. More importantly, we further explore the transfer learning capacity of zero-shot generalisation in unseen environments. We leverage customised dataset comprising 58 testing question-answering pairs, recorded in London, UK, presenting a significant distribution shift from the BDD-X dataset.
Benchmark Settings For all experiments, we train the MLLM using the BDD-X training split. Subsequent evaluation on general explainability and control signal prediction capabilities tests are conducted on the BDD-X test split with the BDD-X training split as the memory database. For the transfer learning experiments, we employ the same foundational model and test it on the Spoken-SAX but the memory database is constructed using the BDD-X training split for zero-shot generalisation.
Implementation Detail For each of the driving videos, we uniformly sample the video to 8 frames and resize it to $224\times 224$ for all frames. For MLLM, we train the model for one and two epochs in the pre-training and fine-tuning stages, respectively. For the embedding projector, we train the model for 300 epochs. Further experiment implementation details are provided in Sec. -B.
[…]
TABLE I: Driving Action Explanation and Justification Evaluation. B4, C, and M represent BLEU4, CIDEr, and METEOR, respectively, for measuring the accuracy and quality of textual driving explanation. The upper part indicates the evaluation on the in-distribution BDD-X test set. The lower part indicates the zero-shot generalisation evaluation on the Spoken-SAX dataset.
Method Is Generalist? Action Justification
B4 $\uparrow$ C $\uparrow$ M $\uparrow$ B4 $\uparrow$ C $\uparrow$ M $\uparrow$
In Distribution Performance (BDD-X)
S2VT [74] $\times$ $30.2$ $179.8$ $27.5$ $6.3$ $53.4$ $11.2$
S2VT++ [74] $\times$ $27.1$ $157.0$ $26.4$ $5.8$ $52.7$ $10.9$
SAA [38] $\times$ $31.8$ $214.8$ $29.1$ 7.1 $66.1$ $12.2$
WAA [38] $\times$ $32.3$ $215.8$ $29.2$ $7.3$ $69.5$ $12.2$
ADAPT [33] $\times$ $34.6$ $247.5$ $30.6$ $11.4$ $102.6$ $15.2$
DriveGPT4 [78] $30.0$ $214.0$ $29.8$ $9.4$ $102.7$ $14.6$
OURS $34.3$ $260.8$ $30.7$ $11.1$ $109.1$ $14.8$
$\Delta$ w.r.t generalist SOTA +$14.3\%$ +$21.9\%$ +$3.0\%$ +$18.1\%$ +$6.2\%$ +$1.4\%$
Zero-shot Generalisation (Spoken-SAX)
ADAPT [33] $\times$ $0.0$ $22.3$ $26.7$ $3.1$ $8.6$ $12.4$
Base w/o ICL $0.0$ $5.7$ $21.6$ $2.1$ $4.7$ $11.4$
OURS w ICL $9.9$ $48.9$ $27.5$ $5.6$ $17.1$ $14.6$
[…]
TABLE II: Control Signals Prediction Accuracy Evaluation on BDD-X dataset.
Method Course Speed
RMSE(degree)$\downarrow$ $A_{0.1}\uparrow$ $A_{0.5}\uparrow$ $A_{1.0}\uparrow$ $A_{5.0}\uparrow$ $A_{10.0}\uparrow$ RMSE(m/s)$\downarrow$ $A_{0.1}\uparrow$ $A_{0.5}\uparrow$ $A_{1.0}\uparrow$ $A_{5.0}\uparrow$ $A_{10.0}\uparrow$
ADAPT [33] $5.87$ $54.49$ $86.39$ $91.06$ $97.36$ $98.20$ $2.68$ $11.77$ $31.79$ $47.48$ $92.75$ $95.87$
DriveGPT4 [78] $4.57$ $69.22$ $79.14$ $84.47$ $95.72$ $96.74$ $1.09$ 56.93 $77.77$ $87.97$ $99.00$ $99.57$
OURS 4.48 74.32 88.69 93.12 98.30 99.10 0.69 $51.12$ 85.54 94.49 99.81 99.91
TABLE III: Ablation Study on Retrieval Strategy and In-Context Learning Phase. Visual and Hybrid Search refers to the use of video embeddings or projected hybrid sensor embeddings, respectively. Training and Inference denote whether in-context learning examples are provided during the corresponding phase.
Retrieval Strategy ICL Phase BDD-X
Visual Search Hybrid Search Training Inference Act. B4 Act. C Jus. B4 Jus. C Speed Err. Course Err.
$29.8$ $222.1$ $5.9$ $58.9$ $0.67$ $5.61$
$0.0$ $0.0$ $0.0$ $0.0$ $N/A$ $N/A$
$0.0$ $0.0$ $0.0$ $0.0$ $N/A$ $N/A$
$31.2$ $222.1$ $7.7$ $83.1$ $0.78$ $5.61$
$34.3$ $260.9$ $11.1$ $109.1$ $0.69$ $4.48$
Figure 4: Qualitative Examples. Comparison between human annotated ground truth (Left) and prediction from our method (Right) from both BDD-X (Top 4) and Spoken-London (Bottom 4). Each video is downsampled to four frames for visualisation. Despite subtle differences in wording, our method provides driving action explanations and justifications that are comprehensible to humans.
Baselines We compare with various driving action description and justification baseline methods such as video-language sequence-to-sequence based recurrent neural network S2VT [74] and visual-attention based convolutional neural network WAA [38]. Also, we compare with the state-of-the-art explainable driving methods, including video transformer-based ADAPT [33] and visual instruction tuned MLLM DriveGPT4 [78] that is further capable of control signal prediction.
IV-B Explainability in Driving Action and Justification
We begin by evaluating the quality and accuracy of explanations and justifications for driving actions separately. As shown in upper part of Tab. I, our method demonstrates comparable performance to the state-of-the-art specialist method ADAPT [33], a characteristic not observed in previous MLLM-based methods. In particular, when compared with DriveGPT4 [78] which also uses MLLM with a similar architecture and number of parameters but incorporates the extra LLaVA-150K dataset [48] for visual instruction tuning, our approach, relying solely on the BDD-X dataset, outperforms it in terms of explainability. This is evidenced by an average performance improvement of 10.8% across all metrics. This underscores the effectiveness of ICL in enhancing the emergent reasoning capabilities of MLLMs.
IV-C Control Signal Prediction
We next evaluate the accuracy of control signal predictions for Course (i.e., turning angle) and Speed. As indicated in Tab. II, our method surpasses others in open-loop control accuracy across various tolerance ranges and in terms of RMSE, significantly outperforming the baseline approach. In particular, when compared to the state-of-the-art DriveGPT4, which also uses the same visual input combined with past control signals for autoregressive prediction, our method stands out by implementing retrieval-augmented ICL examples. This indicates the analogy in the overall reasoning process provided by the ICL examples also contributes to the improvement in numerical control signal prediction.
IV-D Ablation Study on Retrieval Strategy
We perform a more comprehensive ablation study to evaluate the efficacy of our proposed retrieval-augmented in-context learning. We first aim to investigate the similarity metric for retrieval. In particular, we compare the use of visual similarity (i.e., video embeddings only) with hybrid similarity (i.e., hybrid video and control signal projected embedding Sec. III-C). Our empirical findings indicate suboptimal performance when using visual similarity, possibly because it tends to prioritize ICL examples that are most perceptually similar, rather than effectively demonstrating the reasoning process. By fine-tuning the embeddings, we not only harness the potential of heterogeneous multi-modal sensor input but also enable more effective ICL example retrieval.
We also investigate whether to apply ICL examples during training or solely in inference time. As shown in Tab. III, we found that the MLLM is incapable of making reasonable predictions using ICL examples without prior training, regardless of the retrieval strategy chosen. This suggests that the pre-trained MLLM is not equipped to effectively perform zero-shot ICL. We hypothesize that supervised fine-tuning plays a crucial role in enhancing the ICL capabilities of MLLM, necessitating the provision of reasoning demonstrations, which aligns with observations in [18].
IV-E Generalisation Capacity
One of the critical capacities of autonomous systems is to generalise to unseen environments out of its training distribution. However, in the domain of explainable driving, we observe the existing methods are unable to perform such generalisation, which poses challenges for their deployment. As shown in lower part of Tab. I, ADAPT and the base MLLM (i.e., training without ICL) reveal dramatic performance degradation to the in-distribution situation. However, our method leverages ICL learning examples to demonstrate a significant performance boost with a large margin. Note that even though the memory database is constructed with BDD-X, the RA-ICL can still perform generalisation in a zero-shot manner. This is potentially due to the robustness of the hybrid retrieval process, where samples with less distribution shift can still be selected to serve as effective ICL examples.
IV-F Qualitative Demonstration
We also demonstrate a series of quantitative examples comparing the driving action explanation and justification provided by human ground truth and prediction from our method. As shown in Fig. 4, we observe RAG-Driver produces robust intelligible action explanation and justification under different environments (i.e. night time and adversarial weather) with a control signal close to the human driver record. More importantly, in the out-of-distribution setting Spoken-SAX as indicated by the clear visual difference, we observe the prediction also produces human-understandable answers, qualitatively validating the exceptional zero-shot generalisation capacity.
V Limitations and Future Work
This work aims to develop a generalisable explainable driving commentator using a Machine Learning Language Model (MLLM), addressing a significant obstacle that has hindered deployment: the poor generalisation capacity. However, several problems still need to be addressed. For instance, the prevalent issue of hallucination in MLLM, while mitigated, is still observed. We hypothesize that this is due to the limited general video understanding capacity, as the video encoder only processes 8 frames per video. Also, due to the limited access to open-source models and computational cost, we employ a relatively small MLLM with 7 billion parameters, which falls short of some state-of-the-art models (i.e. GPT4V [1], Gemini [70]). We anticipate that the development of a more powerful MLLM backbone with lower computational costs could further enhance the application of MLLM in driving.
VI Conclusion
We propose RAG-Driver, a Multi-Modal Large Language Model with Retrieval-augmented In-context Learning capacity designed for generalisable and explainable end-to-end driving. It exhibits strong capability in providing numerical control signals, along with explanations and justifications for driving actions. More importantly, it shows impressive zero-shot generalisation to unseen environments without the need for additional training.

### v2

@@
 Abstract
-Robots powered by ‘blackbox’ models need to provide human-understandable explanations which we can trust.
-Hence, explainability plays a critical role in trustworthy autonomous decision-making to foster transparency and acceptance among end users, especially in complex autonomous driving.
+We need to trust robots that use often opaque AI methods.
+They need to explain themselves to us, and we need to trust their explanation.
+In this regard, explainability plays a critical role in trustworthy autonomous decision-making to foster transparency and acceptance among end users, especially in complex autonomous driving.
 Recent advancements in Multi-Modal Large Language models (MLLMs) have shown promising potential in enhancing the explainability as a driving agent by producing control predictions along with natural language explanations.
@@
 By grounding in retrieved expert demonstration, we empirically validate that RAG-Driver achieves state-of-the-art performance in producing driving action explanations, justifications, and control signal prediction.
-More importantly, it exhibits exceptional zero-shot generalisation capabilities to unseen environments without further training endeavours [^1] Project Page: https://yuanjianhao508.github.io/RAG-Driver.
+More importantly, it exhibits exceptional zero-shot generalisation capabilities to unseen environments without further training endeavours.
 Index Terms:
@@
 I Introduction
-Driven by the emerging development of deep learning, autonomous driving has observed a paradigm shift from rules-based decision systems [66, 21] to data-driven learning-based approaches [28, 6, 36].
-However, this comes at the cost of transparency in decision-making, especially for end-to-end autonomous driving systems which are considered black-box in nature [13].
-Thus, in addition to precision in action control, explanation provision is key in ensuring trustworthy decision-making to reconcile the system’s decisions with end-user expectations to foster confidence and acceptance [79, 8, 57] in dynamic driving environments.
-Traditional approaches have mainly relied on attention visualisation [5, 7, 55] as a proxy to rationalise the decisions of the black-box systems or auxiliary intermediate tasks such as semantic segmentation [25, 32], object detection [16, 31], and affordance prediction [68, 45] provide meaningful intermediate representation for decision-making.
+Driven by the emerging development of deep learning, autonomous driving has observed a paradigm shift from rules-based decision systems [schwarting2018planning, faisal2019understanding] to data-driven learning-based approaches [Hu_2023_CVPR, bojarski2016end, kendall2019learning].
+However, this comes at the cost of transparency in decision-making, especially for end-to-end autonomous driving systems which are considered black-box in nature [chen2023end].
+Thus, in addition to precision in action control, explanation provision is key in ensuring trustworthy decision-making to reconcile the system’s decisions with end-user expectations to foster confidence and acceptance [yang2023survey, bonnefon2016social] in dynamic driving environments.
+Traditional approaches have mainly relied on attention visualisation [bojarski2016visualbackprop, bojarski2017explaining, moriVisualExpl] as a proxy to rationalise the decisions of the black-box systems or auxiliary intermediate tasks such as semantic segmentation [hawke2020urban, jain2023ground], object detection [chitta2022transfuser, jaeger2023hidden], and affordance prediction [shao2023safety, li2018rethinking] provide meaningful intermediate representation for decision-making.
 However, these methods do not engage end-users in the dialogue as they are one-directional and not readily comprehensible by the general users for the purpose of fostering trust and confidence.
-An alternative promising approach is the integration of natural language explanations [38, 33, 54], in particular through Multi-Modal Large Language Models (MLLMs) [1, 70].
+An alternative promising approach is the integration of natural language explanations [kim2018bddx, jin2023adapt, marcu2023lingoqa], in particular through Multi-Modal Large Language Models (MLLMs) [achiam2023gpt, team2023gemini].
 These models, pre-trained on extensive web-scale datasets, demonstrate remarkable reasoning capacity, enabling the transformation of complex vehicular decision-making processes into more understandable narrative formats, thereby offering a new layer of explainability to conventional systems.
-While several early attempts have demonstrated the potential of MLLMs as general explainable driving agents [78, 76, 51], these methods fall short of human-level understanding.
+While several early attempts have demonstrated the potential of MLLMs as general explainable driving agents [xu2023drivegpt4, wang2023drivemlm, ma2023dolphins], these methods fall short of human-level understanding.
 One of the limitations is their failure to generalise to unseen environments.
-A primary obstacle is the lack of high-quality annotated data [56], coupled with the significant domain shift across various datasets [23], which hinders the models’ generalisation capacity to novel environments outside of the training data distribution.
-Another critical challenge is the prohibitively expensive training requirement and the unsolved problem of catastrophic forgetting [39], which make re-training or fine-tuning impractical solutions due to the immense computational demands and severe performance degradation.
+A primary obstacle is the lack of high-quality annotated data [nie2023reason2drive], coupled with the significant domain shift across various datasets [domainbed], which hinders the models’ generalisation capacity to novel environments outside of the training data distribution.
+Another critical challenge is the prohibitively expensive training requirement and the unsolved problem of catastrophic forgetting [kirkpatrick2017overcoming], which make re-training or fine-tuning impractical solutions due to the immense computational demands and severe performance degradation.
 Consequently, this further limits the models’ generalisability after deployment, as they struggle to effectively utilise new data in constantly evolving environments and driving scenarios.
@@
 As illustrated in Fig.1, it outputs natural language texts corresponding to (1) the driving action and (2) justification of that driving action along with (3) numerical control signals based on driving videos.
-The natural language texts are aligned with control signals during in-context learning [10] to enable faithful introspective explanation provision.
+The natural language texts are aligned with control signals during in-context learning [brown2020language] to enable faithful introspective explanation provision.
 The novelty of RAG-Driver is the integration of retrieval-augmented in-context learning (RA-ICL) mechanisms that significantly improve generalisation performance in unseen driving environments.
@@
   2. 2.
-    Achieving state-of-the-art introspective driving explanation performance on the standard benchmark BDD-X [38].
+    Achieving state-of-the-art introspective driving explanation performance on the standard benchmark BDD-X [kim2018bddx].
   3. 3.
-    Demonstrating exceptional zero-shot generalisation capacity to unseen scenarios without training effort through a customised dataset Spoken-SAX featuring video sequences narrated by a professional driving instructor.
+    Demonstrating exceptional zero-shot generalisation capacity to unseen scenarios without training effort through a customised dataset Spoken-London featuring video sequences narrated by a professional driving instructor.
 II Related Work
 II-A Explainable End-to-End Autonomous Driving
-End-to-end learned driving [13] maps directly from raw sensor input to vehicle control signals.
-This data-driven, joint optimisation of perception, prediction, and planning can be simple and efficient [28].
-In this area, various learning-based approaches, including behavior cloning [6, 12, 61, 83, 60], inverse optimal control [81, 65, 75], and reinforcement learning [36, 71, 11, 44] are promising.
-A key area of focus in this area is explainability [80], which is crucial for improving transparency and building trust towards wider public acceptance of autonomous systems [54, 22, 57].
-One line of works leverages attention visualisation – either to directly identify salient regions of input images that are important for driving decision-making [37, 5, 7] or to assist feature aggregation for downstream motion planning tasks [55, 15, 77, 63].
-Another line of work uses intermediate auxiliary tasks such as semantic segmentation [25, 32], object detection [16, 31], and affordance prediction [68, 45] which help to decoder latent representations to human-understandable representation.
+End-to-end learned driving [chen2023end] maps directly from raw sensor input to vehicle control signals.
+This data-driven, joint optimisation of perception, prediction, and planning can be simple and efficient [Hu_2023_CVPR].
+In this area, various learning-based approaches, including behavior cloning [bojarski2016end, chen2020learning, pomerleau1988alvinn, zhang2022learning, peng2023learning], inverse optimal control [zeng2019end, sadat2020perceive, wang2021end], and reinforcement learning [kendall2019learning, toromanoff2020end, chekroun2023gri, li2022efficient] are promising.
+A key area of focus in this area is explainability [zablocki2022explainability], which is crucial for improving transparency and building trust towards wider public acceptance of autonomous systems [marcu2023lingoqa, gadd2020sense, OmeizaExplainSurvey2022].
+One line of works leverages attention visualisation – either to directly identify salient regions of input images that are important for driving decision-making [kim2017interpretable, bojarski2016visualbackprop, bojarski2017explaining] or to assist feature aggregation for downstream motion planning tasks [moriVisualExpl, chitta2021neat, xiao2023scaling, renz2022plant].
+Another line of work uses intermediate auxiliary tasks such as semantic segmentation [hawke2020urban, jain2023ground], object detection [chitta2022transfuser, jaeger2023hidden], and affordance prediction [shao2023safety, li2018rethinking] which help to decoder latent representations to human-understandable representation.
 While these methods provide explainable mechanisms by associating decision-making processes with semantic or visual representations, they are not readily comprehensible by the general users for the purpose of fostering trust and confidence.
 Alternatively, recent research shows promise in utilising natural language explanation.
-Several works develop specialist explainers [38, 41] using attention alignment in visual input and textual generation for grounded driving action explanation.
-ADAPT [33] uses a vision-language transformer with separate decoder for caption generation alongside control signal prediction.
+Several works develop specialist explainers [kim2018bddx, kuhn2023textual] using attention alignment in visual input and textual generation for grounded driving action explanation.
+ADAPT [jin2023adapt] uses a vision-language transformer with separate decoder for caption generation alongside control signal prediction.
 More recently, several works explore the potential of Multi-modal Large Language Models (MLLMs).
-Works such as, DriveGPT4 [78], Lingo [54], and DrivingMLM [76] have shown promising potential in general question-answering for driving and action planning.
+Works such as, DriveGPT4 [xu2023drivegpt4], Lingo [marcu2023lingoqa], and DrivingMLM [wang2023drivemlm] have shown promising potential in general question-answering for driving and action planning.
 However, a common obstacle in both specialist and MLLM-based generalist models is the scarcity of data due to expensive annotation costs and significant domain gaps between different datasets, making the development of a robust and generalisable model an extremely challenging task.
@@
 II-B Multi-Modal Large Language Model
-Recent advancements in Large Language Models (LLMs) have paved the way for the emergence of Multi-modal Large Language Models (MLLMs) [1, 70].
+Recent advancements in Large Language Models (LLMs) have paved the way for the emergence of Multi-modal Large Language Models (MLLMs) [achiam2023gpt, team2023gemini].
 Benefiting from scalable transformer-based architecture and web-scale training data, these models have demonstrated notable capabilities in general-purpose visual understanding tasks.
 One line of work focuses on modality fusion in the latent space, offering a scalable, end-to-end solution for MLLMs.
-For instance, Flamingo [3] and BLIP2 [43] fuse visual tokens to frozen LLM through gated attention and query transformers, respectively.
-LLaVA [48] and MiniGPT4 [86] use simple Multi-Layer Perceptron (MLP) with visual instruction tuning to align the pre-trained image encoder to LLM.
-Most relevant to us is the line of work focusing on video-language models such as Video-LLaVA [46] and Video-LLaMA [82], which integrate pre-trained video encoders into LLMs using strategies similar to those in image-based models.
-With remarkable perception and reasoning capacity, MLLMs reveal promising potential in various robotics tasks such as reasoning [20, 2, 30] and planning [62, 9, 47, 35].
+For instance, Flamingo [alayrac2022flamingo] and BLIP2 [li2023blip] fuse visual tokens to frozen LLM through gated attention and query transformers, respectively.
+LLaVA [liu2023llava] and MiniGPT4 [zhu2023minigpt] use simple Multi-Layer Perceptron (MLP) with visual instruction tuning to align the pre-trained image encoder to LLM.
+Most relevant to us is the line of work focusing on video-language models such as Video-LLaVA [lin2023video] and Video-LLaMA [damonlpsg2023videollama], which integrate pre-trained video encoders into LLMs using strategies similar to those in image-based models.
+With remarkable perception and reasoning capacity, MLLMs reveal promising potential in various robotics tasks such as reasoning [driess2023palm, ahn2022can, huang2023instruct2act] and planning [rana2023sayplan, brohan2023rt, lin2023text2motion].
 The most similar to us is the idea that engages a generalist foundation model for end-to-end embodied agent.
-PaLM-e [20] injects images, state estimates, and other sensor modalities to LLM and autoregressively produces natural language commands.
-RT-2 [9] and RT-X [58] fine-tune on image and low-level robot control signal pairs to perform end-to-end robot control.
-Specifically in driving, numerous approaches leverage language-only LLM for decision-making [79] and then augment it with external perception module feedback [53, 24, 14], designed chain-of-thought reasoning template [53, 52] or downstream planner [67, 69, 14] to form a system-level driving agent.
+PaLM-e [driess2023palm] injects images, state estimates, and other sensor modalities to LLM and autoregressively produces natural language commands.
+RT-2 [brohan2023rt] and RT-X [padalkar2023open] fine-tune on image and low-level robot control signal pairs to perform end-to-end robot control.
+Specifically in driving, numerous approaches leverage language-only LLM for decision-making [yang2023survey] and then augment it with external perception module feedback [mao2023language, han2024dme, chen2023driving], designed chain-of-thought reasoning template [mao2023language, ma2023lampilot] or downstream planner [sha2023languagempc, sharan2023llm, chen2023driving] to form a system-level driving agent.
 Another more relevant line of work is end-to-end driving agents.
-DriveGPT4 [78] leverages a finetuned video-language model Valley [50] on driving specific visual instruction tuning based on BDD-X [38].
-Dolphins [51] further use the designed Grounded chain-of-thought to enhance reasoning capacity.
-DrivingMLM [76] and Reason2Drive [56] scale up the driving visual instruction tuning dataset through simulator-based data engine and annotation of existing large-scale datasets, respectively.
+DriveGPT4 [xu2023drivegpt4] leverages a finetuned video-language model Valley [luo2023valley] on driving specific visual instruction tuning based on BDD-X [kim2018bddx].
+Dolphins [ma2023dolphins] further use the designed Grounded chain-of-thought to enhance reasoning capacity.
+DrivingMLM [wang2023drivemlm] and Reason2Drive [nie2023reason2drive] scale up the driving visual instruction tuning dataset through simulator-based data engine and annotation of existing large-scale datasets, respectively.
 While these approaches have demonstrated the potentials of MLLM, the prohibitively expensive training cost and the unsolved problem of catastrophic forgetting, which makes re-training or fine-tuning post-deployment challenging, further limit their generalisation capacity to unseen driving environments.
@@
 II-C In-Context Learning and Retrieval-Augmented Generation
-While LLMs demonstrate strong generative and reasoning capacity, there are still several issues associated with their output, such as hallucination [29], and slow knowledge updates [34].
-In-context Learning (ICL) [10, 18] has emerged as a promising approach in LLM inference, potentially addressing several of these issues.
+While LLMs demonstrate strong generative and reasoning capacity, there are still several issues associated with their output, such as hallucination [huang2023survey], and slow knowledge updates [kandpal2023large].
+In-context Learning (ICL) [brown2020language, dong2022survey] has emerged as a promising approach in LLM inference, potentially addressing several of these issues.
 This paradigm involves providing a test query alongside a few demonstration examples as contextual information.
-The LLM then generates an output for the test instance based on analogies drawn from context, without any updates to its parameters [64].
+The LLM then generates an output for the test instance based on analogies drawn from context, without any updates to its parameters [rubin2021learning].
 While ICL has been observed to enhance generalisability in various Natural Language Processing (NLP) tasks, its application in multi-modal contexts remains less explored, potentially due to the challenges associated with curating structured high-quality multi-modal ICL datasets.
-Retrieval-Augmented Generation (RAG) [42] is another important inference paradigm for LLM.
+Retrieval-Augmented Generation (RAG) [lewis2020retrieval] is another important inference paradigm for LLM.
 It provides an external knowledge database to augment the compressed model knowledge within LLM in inference by dynamically retrieving relevant information pieces as contextual information.
@@
 III-A Multi-modal Large Language Model Architecture
-Following the successful MLLM paradigm of Video-LLaVA [46], we align visual and language embedding through visual instruction tuning.
+Following the successful MLLM paradigm of Video-LLaVA [lin2023video], we align visual and language embedding through visual instruction tuning.
 We leverage a pre-trained video encoder and LLM and then inject the video embedding into LLM through a MLP projector to build a fully differentiable MLLM.
-Video Encoder We adopt the pre-trained LanguageBind video encoder [85] as our frozen visual backbone $f_{v}$, which is based on a ViT-B/32 vision transformer [19].
+Video Encoder We adopt the pre-trained LanguageBind video encoder [zhu2023languagebind] as our frozen visual backbone $f_{v}$, which is based on a ViT-B/32 vision transformer [dosovitskiy2020image].
 As shown in Fig.
@@
 The video encoder is pre-trained with video-language contrastive learning (i.e.
-CLIP4clip [49]) without further fine-tuning.
+CLIP4clip [Luo2021CLIP4Clip]) without further fine-tuning.
 Figure 2: Video Encoder architecture.
 Video is first split into $k\times 32\times 32$ patches concatenated in time, where these patches are linear projected to video embedding.
-Then, the model is trained with video-language contrastive learning (CLIP4clip) [49] to obtain language-align video representation.
+Then, the model is trained with video-language contrastive learning (CLIP4clip) [Luo2021CLIP4Clip] to obtain language-align video representation.
 Cross-Modality Projector We then leverage a two-layer MLP to project and align the encoded video embedding $\bm{z}_{vo}$ with language token embeddings $\bm{z}_{v}\in\mathbb{R}^{2048\times 4096}$.
 $f_{p}(\bm{z}_{vo})=\text{GELU}\left(W_{2}\cdot\text{GELU}\left(W_{1}\cdot\bm{z}_{vo}\right)\right)$ (1)
-In particular, the projector $f_{p}$ takes form in Eq. 1, where we use GELU [26] as an activation function. We train the projector with a two-stage training strategy as detailed in Sec. III-B.
+In particular, the projector $f_{p}$ takes form in Eq.
+1, where we use GELU [hendrycks2016gaussian] as an activation function.
+We train the projector with a two-stage training strategy as detailed in Sec.
+III-B.
 Large Language Model Backbone Finally, the LLM takes in both aligned video embedding $\bm{z}_{v}$ and language embedding $\bm{z}_{t}$ of textual context information and task instruction to predict both textual action explanation and numerical control signal.
-We adopt Vicuna 1.5 7B [84], which is instruction-tuned based on LLaMA2 [72] as our LLM backbone.
+We adopt Vicuna 1.5 7B [zheng2023judging], which is instruction-tuned based on LLaMA2 [touvron2023llama2] as our LLM backbone.
 For decoder-only LLM conditioned on the multi-modal contextual prefix $\bm{z}_{1:n}=[\bm{z}_{v},\bm{z}_{t}]$ of length $N$, the joint probability of output $\bm{x}_{n+1:L}$ as in Eq.
@@
 III-B Training Strategy
-Following the visual instruction tuning paradigm [48, 46], we employ a two-stage training strategy to progressively enable cross-modality alignment and multi-task driving capacity.
+Following the visual instruction tuning paradigm [liu2023llava, lin2023video], we employ a two-stage training strategy to progressively enable cross-modality alignment and multi-task driving capacity.
 In both stages, we leverage the same next-token prediction cross-entropy loss as in Eq.
@@
 Pre-training In the first pre-training stage, we only train the cross-modality projector while freezing the visual encoder and LLM.
-We use a subset VIDAL-10M [85] which contains 3 million video-caption pairs.
+We use a subset VIDAL-10M [zhu2023languagebind] which contains 3 million video-caption pairs.
 This achieves an alignment between visual and language features by projecting the pre-trained video embeddings into language tokens understandable by LLM.
-Supervised In-context Instruction Tuning While state-of-the-art LLM exhibits zero-shot ICL capability, several works [18] and our ablation study Sec.
+Supervised In-context Instruction Tuning While state-of-the-art LLM exhibits zero-shot ICL capability, several works [dong2022survey] and our ablation study Sec.
 IV-D have shown further improvement when specifically trained on curated ICL demonstrations.
-In particular, we construct a multi-modal instruction tuning dataset with structured ICL examples based on the BDD-X dataset [38] resulting in 16K video question and answer pairs.
+In particular, we construct a multi-modal instruction tuning dataset with structured ICL examples based on the BDD-X dataset [kim2018bddx] resulting in 16K video question and answer pairs.
 As shown in Fig.
@@
 Retrieval Mechanism To perform the retrieval, we first leverage a lightweight MLP projector of the same structure as in Eq.
-1 to project the heterogeneous video and control signal embedding into the same hybrid embedding $\bm{s}\in\mathbb{R}^{1024}$ through metric learning [27].
+1 to project the heterogeneous video and control signal embedding into the same hybrid embedding $\bm{s}\in\mathbb{R}^{1024}$ through metric learning [hoffer2015deep].
 In particular, we adopt a triplet loss with Euclidean distance as shown in Eq.
@@
 3.
-Retrieval Augmented In-Context Learning (RA-ICL) To perform the RA-ICL, we prefix retrieved samples before the current query, facilitating an implicit gradient descent through the meta-optimiser of LLM, as proved in [17].
+Retrieval Augmented In-Context Learning (RA-ICL) To perform the RA-ICL, we prefix retrieved samples before the current query, facilitating an implicit gradient descent through the meta-optimiser of LLM, as proved in [dai2022can].
 This approach is also applicable to MLLM with architecture specified in Sec.
@@
 $W_{Q,K,V}\in\mathbb{R}^{d^{i}\times d^{o}}$ is the linear transformation on query, key, and value in attention block, respectively.
-Now, in [17] in-context learning was shown to effectively achieve a meta-optimisation to implicitly update the model with an estimated meta-gradient.
-We provide in our work a new, alternative derivation of this, supported by more recent work in [40] on the following softmax-free linear attention expressions
+Now, in [dai2022can] in-context learning was shown to effectively achieve a meta-optimisation to implicitly update the model with an estimated meta-gradient.
+We provide in our work a new, alternative derivation of this, supported by more recent work in [koohpayegani2024sima] on the following softmax-free linear attention expressions
 $\mathcal{F}_{\mathrm{ICL}}(\bm{z}_{1:n})=W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\left(W_{K}\left[\bm{z}_{icl};\bm{z}_{q}\right]\right)^{T}W_{Q}\bm{z}_{1:n}$ (7)
@@
 (2) Control Signal Prediction.
-We conduct experiments with the BDD-X [38] dataset, which is a widely adopted benchmark in explainable driving, comprising 77-hour videos across the US under different road and weather conditions.
+We conduct experiments with the BDD-X [kim2018bddx] dataset, which is a widely adopted benchmark in explainable driving, comprising 77-hour videos across the US under different road and weather conditions.
 We customize the format as shown in Fig.
@@
 Subsequent evaluation on general explainability and control signal prediction capabilities tests are conducted on the BDD-X test split with the BDD-X training split as the memory database.
-For the transfer learning experiments, we employ the same foundational model and test it on the Spoken-SAX but the memory database is constructed using the BDD-X training split for zero-shot generalisation.
+For the transfer learning experiments, we employ the same foundational model and test it on the Spoken-London but the memory database is constructed using the BDD-X training split for zero-shot generalisation.
 Implementation Detail For each of the driving videos, we uniformly sample the video to 8 frames and resize it to $224\times 224$ for all frames.
@@
 In Distribution Performance (BDD-X)
-S2VT [74] $\times$ $30.2$ $179.8$ $27.5$ $6.3$ $53.4$ $11.2$
-S2VT++ [74] $\times$ $27.1$ $157.0$ $26.4$ $5.8$ $52.7$ $10.9$
-SAA [38] $\times$ $31.8$ $214.8$ $29.1$ 7.1 $66.1$ $12.2$
-WAA [38] $\times$ $32.3$ $215.8$ $29.2$ $7.3$ $69.5$ $12.2$
-ADAPT [33] $\times$ $34.6$ $247.5$ $30.6$ $11.4$ $102.6$ $15.2$
-DriveGPT4 [78] $30.0$ $214.0$ $29.8$ $9.4$ $102.7$ $14.6$
+S2VT [venugopalan2015sequence] $\times$ $30.2$ $179.8$ $27.5$ $6.3$ $53.4$ $11.2$
+S2VT++ [venugopalan2015sequence] $\times$ $27.1$ $157.0$ $26.4$ $5.8$ $52.7$ $10.9$
+SAA [kim2018bddx] $\times$ $31.8$ $214.8$ $29.1$ 7.1 $66.1$ $12.2$
+WAA [kim2018bddx] $\times$ $32.3$ $215.8$ $29.2$ $7.3$ $69.5$ $12.2$
+ADAPT [jin2023adapt] $\times$ $34.6$ $247.5$ $30.6$ $11.4$ $102.6$ $15.2$
+DriveGPT4 [xu2023drivegpt4] $30.0$ $214.0$ $29.8$ $9.4$ $102.7$ $14.6$
 OURS $34.3$ $260.8$ $30.7$ $11.1$ $109.1$ $14.8$
 $\Delta$ w.r.t generalist SOTA +$14.3\%$ +$21.9\%$ +$3.0\%$ +$18.1\%$ +$6.2\%$ +$1.4\%$
-Zero-shot Generalisation (Spoken-SAX)
-ADAPT [33] $\times$ $0.0$ $22.3$ $26.7$ $3.1$ $8.6$ $12.4$
+Zero-shot Generalisation (Spoken-SAX) B4 $\uparrow$ C $\uparrow$ M $\uparrow$ B3 $\uparrow$ C $\uparrow$ M $\uparrow$
+ADAPT [jin2023adapt] $\times$ $0.0$ $22.3$ $26.7$ $3.1$ $8.6$ $12.4$
 Base w/o ICL $0.0$ $5.7$ $21.6$ $2.1$ $4.7$ $11.4$
@@
 RMSE(degree)$\downarrow$ $A_{0.1}\uparrow$ $A_{0.5}\uparrow$ $A_{1.0}\uparrow$ $A_{5.0}\uparrow$ $A_{10.0}\uparrow$ RMSE(m/s)$\downarrow$ $A_{0.1}\uparrow$ $A_{0.5}\uparrow$ $A_{1.0}\uparrow$ $A_{5.0}\uparrow$ $A_{10.0}\uparrow$
-ADAPT [33] $5.87$ $54.49$ $86.39$ $91.06$ $97.36$ $98.20$ $2.68$ $11.77$ $31.79$ $47.48$ $92.75$ $95.87$
-DriveGPT4 [78] $4.57$ $69.22$ $79.14$ $84.47$ $95.72$ $96.74$ $1.09$ 56.93 $77.77$ $87.97$ $99.00$ $99.57$
+ADAPT [jin2023adapt] $5.87$ $54.49$ $86.39$ $91.06$ $97.36$ $98.20$ $2.68$ $11.77$ $31.79$ $47.48$ $92.75$ $95.87$
+DriveGPT4 [xu2023drivegpt4] $4.57$ $69.22$ $79.14$ $84.47$ $95.72$ $96.74$ $1.09$ 56.93 $77.77$ $87.97$ $99.00$ $99.57$
 OURS 4.48 74.32 88.69 93.12 98.30 99.10 0.69 $51.12$ 85.54 94.49 99.81 99.91
@@
 $34.3$ $260.9$ $11.1$ $109.1$ $0.69$ $4.48$
-Figure 4: Qualitative Examples.
-Comparison between human annotated ground truth (Left) and prediction from our method (Right) from both BDD-X (Top 4) and Spoken-London (Bottom 4).
-Each video is downsampled to four frames for visualisation.
-Despite subtle differences in wording, our method provides driving action explanations and justifications that are comprehensible to humans.
-Baselines We compare with various driving action description and justification baseline methods such as video-language sequence-to-sequence based recurrent neural network S2VT [74] and visual-attention based convolutional neural network WAA [38].
-Also, we compare with the state-of-the-art explainable driving methods, including video transformer-based ADAPT [33] and visual instruction tuned MLLM DriveGPT4 [78] that is further capable of control signal prediction.
+[…]
+Baselines We compare with various driving action description and justification baseline methods such as video-language sequence-to-sequence based recurrent neural network S2VT [venugopalan2015sequence] and visual-attention based convolutional neural network WAA [kim2018bddx].
+Also, we compare with the state-of-the-art explainable driving methods, including video transformer-based ADAPT [jin2023adapt] and visual instruction tuned MLLM DriveGPT4 [xu2023drivegpt4] that is further capable of control signal prediction.
 IV-B Explainability in Driving Action and Justification
@@
 As shown in upper part of Tab.
-I, our method demonstrates comparable performance to the state-of-the-art specialist method ADAPT [33], a characteristic not observed in previous MLLM-based methods.
-In particular, when compared with DriveGPT4 [78] which also uses MLLM with a similar architecture and number of parameters but incorporates the extra LLaVA-150K dataset [48] for visual instruction tuning, our approach, relying solely on the BDD-X dataset, outperforms it in terms of explainability.
+I, our method demonstrates comparable performance to the state-of-the-art specialist method ADAPT [jin2023adapt], a characteristic not observed in previous MLLM-based methods.
+In particular, when compared with DriveGPT4 [xu2023drivegpt4] which also uses MLLM with a similar architecture and number of parameters but incorporates the extra LLaVA-150K dataset [liu2023llava] for visual instruction tuning, our approach, relying solely on the BDD-X dataset, outperforms it in terms of explainability.
 This is evidenced by an average performance improvement of 10.8% across all metrics.
@@
 This suggests that the pre-trained MLLM is not equipped to effectively perform zero-shot ICL.
-We hypothesize that supervised fine-tuning plays a crucial role in enhancing the ICL capabilities of MLLM, necessitating the provision of reasoning demonstrations, which aligns with observations in [18].
+We hypothesize that supervised fine-tuning plays a crucial role in enhancing the ICL capabilities of MLLM, necessitating the provision of reasoning demonstrations, which aligns with observations in [dong2022survey].
 IV-E Generalisation Capacity
@@
 night time and adversarial weather) with a control signal close to the human driver record.
-More importantly, in the out-of-distribution setting Spoken-SAX as indicated by the clear visual difference, we observe the prediction also produces human-understandable answers, qualitatively validating the exceptional zero-shot generalisation capacity.
+More importantly, in the out-of-distribution setting Spoken-London as indicated by the clear visual difference, we observe the prediction also produces human-understandable answers, qualitatively validating the exceptional zero-shot generalisation capacity.
 V Limitations and Future Work
@@
 Also, due to the limited access to open-source models and computational cost, we employ a relatively small MLLM with 7 billion parameters, which falls short of some state-of-the-art models (i.e.
-GPT4V [1], Gemini [70]).
+GPT4V [achiam2023gpt], Gemini [team2023gemini]).
 We anticipate that the development of a more powerful MLLM backbone with lower computational costs could further enhance the application of MLLM in driving.
@@
 More importantly, it shows impressive zero-shot generalisation to unseen environments without the need for additional training.
+-A Discussion and Limitation
+Scale of MLLM While our Multi-modal Large Language Model (MLLM) has exhibited impressive capabilities in visual reasoning and planning for driving tasks, it’s worth noting that it comprises only 7 billion parameters.
+This size is relatively modest when compared to more well-known models such as GPT4-V [achiam2023gpt] and Gemini [team2023gemini], which boast a significantly larger parameter count and demonstrate superior, near-human levels of visual understanding and reasoning.
+In various related fields such as visual question answering, problem-solving, and interactive dialogue, a clear trend has been observed by researchers: the scale of the model’s parameters and the breadth of training data sources are pivotal.
+Performance improvements are typically seen in tandem with model scaling.
+Based on this trend, we expect similar advancements in the realm of driving applications.
+A larger model could further enhance the capabilities of MLLM in driving scenarios.
+Number of In-Context Learning Examples During training and inference, we provide 2 ICL examples for each query.
+This is due to the limitation of the context window size of the LLM backbone of 2048.
+With the recent improvement in LLM context window size, we expect to see a more flexible adoption of ICL examples.
+-B Training Details
+Embedding Projector We leverage a three-layer MLP as the embedding projector to fuse the video $1\times 1024$ and control signal $1\times 28$ embedding into a hybrid embedding.
+The architecture of the lightweight projector is a four-layer MLP with GELU activation.
+It has an input dimension of $1052$ and an output dimension of $1024$.
+The margin used in triplet loss is 0.5.
+We train the model for 200 epochs with a learning rate of $51e-$ 5 with Adam optimiser.
+MLLM Backbone We use a learning rate of $52e-$ 5 with a cosine scheduler.
+We use a batch size of 4 with a gradient accumulation step of 2 on 8 A100 GPUs, which leads to an effective training batch size of 128.
+We use a warm-up strategy in the first 5 epochs with a warm-up ratio of 0.03.
+We train the model for 2 epochs.
+-C Baseline Details
+In our comparison, we evaluate several baseline methods.
+The first, S2VT [venugopalan2015sequence], utilises an end-to-end sequence-to-sequence model with Long Short-Term Memory (LSTM) networks.
+It is trained on paired video-sentence data, linking video frame sequences to corresponding word sequences, enabling it to generate descriptive captions for video events.
+The second method, WAA [kim2018bddx], employs a visual attention model that trains a convolutional network from images to vehicle control commands.
+This method focuses on identifying influential image regions through the controller’s attention and uses an attention-based video-to-text model to produce textual explanations aligned with the controller’s attention maps, grounding the explanations in relevant scene parts.
+The third approach, ADAPT [jin2023adapt], is a transformer-based method that leverages a multi-task joint training framework.
+It aligns driving action captioning with control signal prediction tasks.
+Finally, DriveGPT4 [xu2023drivegpt4], trained using LLaVA [liu2023llava] to generate a visual instruction tuning dataset derived from the BDD-X dataset [kim2018bddx], processes multimodal input data and is capable of generating text responses while predicting control signals, fine-tuned with the assistance of ChatGPT on the new dataset.
+-D Linear Layer Parameter Update Derivations
+Consider a forward-pass through a linear layer $\mathcal{F}$ after its weights $W_{0}$ have been updated by $\Delta{}W$
+$\mathbf{y}=\mathcal{F}(\mathbf{x})=(W_{0}+\Delta{}W)\mathbf{x}=W_{0}\mathbf{x}+\Delta{}W\mathbf{x}$
+The weight update itself is expressed as
+$\Delta{}W=\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\Rightarrow\Delta{}W\mathbf{x}=\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$
+where $\mathbf{x}_{i},\mathbf{y}_{i}$ are the input and output to the layer that resulted in the weight update.
+Now, if we in fact have optimised over a mini-batch of input-outputs $\mathbf{x}_{i},\mathbf{y}_{i}$ we have
+$\Delta{}W=\sum_{i}\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\Rightarrow\Delta{}W\mathbf{x}=\sum_{i}\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$
+Therefore we have a weighted sum of dot products, which is akin to an attention mechanism.
+Indeed, from Eq.
+6 we can apply softmax-free linear attention expressions such as in [koohpayegani2024sima] for
+$W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\mathrm{softmax}\left(\frac{\left(W_{K}\bm{z}_{1:n}\right)^{T}(W_{Q}\bm{z}_{1:n})}{\sqrt{d^{i}}}\right)$
+$\rightarrow W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\left(W_{K}\left[\bm{z}_{icl};\bm{z}_{q}\right]\right)^{T}W_{Q}\bm{z}_{1:n}$
+Multiplying the linear attention matrices through the stacked in-context and query embeddings we have
+$W_{V}\bm{z}_{icl}(W_{K}\bm{z}_{icl})^{T}W_{Q}\bm{z}_{1:n}+W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T}W_{Q}\bm{z}_{1:n}$
+Now take out a common factor of $W_{Q}\bm{z}_{1:n}$ for
+$(W_{V}\bm{z}_{icl}(W_{K}\bm{z}_{icl})^{T}+W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T})W_{Q}\bm{z}_{1:n}$
+and put $\Delta{}W_{ICL}=W_{V}\bm{z}_{icl}(W_{K}\bm{z}_{icl})^{T}$ and $W_{ZSL}=W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T}$ as the terms both pre-multiplying $W_{Q}\bm{z}_{1:n}$.
+Note that $W_{ZSL}$ is independent of the in-context terms (depending only on the query).
+Now, in $\Delta{}W_{ICL}$ we in fact have a set of in-context retrieved samples $\bm{z}_{icl,i}$ such that
+$W_{V}\left[\bm{z}_{icl,0},\bm{z}_{icl,1},\ldots\right](W_{K}\left[\bm{z}_{icl,0},\bm{z}_{icl,1},\ldots\right])^{T}$
+$\rightarrow\Delta{}W_{ICL}=\sum_{i}~{}W_{V}\bm{z}_{icl,i}(W_{K}\bm{z}_{icl,i})^{T}$
+Finally, by inspection of similar dot-product expressions $W_{V}\bm{z}_{icl,i}(W_{K}\bm{z}_{icl,i})^{T}\leftrightarrow\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$ we note that we match the form for the linear layer above
+$(W_{ZSL}+\Delta{}W_{ICL})W_{Q}\bm{z}_{1:n}\leftrightarrow(W+\Delta{}W)\mathbf{x}$
+This can therefore be interpreted to say that the output of the attention is adjusted in a meta-optimal way to conform to the samples provided as input context, much like gradient descent on the linear layer would adjust that layer to conform to the mini-batch training data, but crucially in the case of RAG-Driver, without supervision.